A survey of 100 similar-sized hospitals revealed a mean daily census in the pediatrics service of 27with a standard deviation of 6.5. Do these data provide sufficient evidence to indicate that thepopulation mean is greater than 25? Let a: 0,05
1
Expert's answer
2019-06-27T13:34:55-0400

Solution.  We define null and alternative hypothesis:


"H_0: \mu \leq 25,"

"H_a: \mu >25."

Next we find a critical value. We will use a right-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

The significance level is "\alpha = 0.05", and the critical value for a right-tailed test is 

"t_c = t(\alpha, n-1)= t(0.05, 99) = 1.66."

The rejection region for this right-tailed test is "t > 1.66" .

The t-statistic is computed as follows:


"t = \frac{\bar{x} - \mu}{s}\cdot \sqrt{n} = \frac{27 - 25}{6.5}\cdot \sqrt{100} = \frac{2}{6.5}\cdot 10 = 3.077."

Since it is observed that "t = 3.077 > t_c = 1.66", it is then concluded that the null hypothesis is rejected.

Conclusion:

It is concluded that the null hypothesis "H_0"  is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is greater than 25, at the 0.05 significance level.